OpenStudy (blackops2luvr):

-7x + y = -12 7x + 5y = -18

OpenStudy (anonymous):

you put the -7x to the other side of the equal sign then you'll get y = -12 +7x then you plug the -12 + 7x into y, the equation at the bottom an put it in () it would look like this 7x + 5(-12 +7x) = -12

OpenStudy (anonymous):

Do you know how to solve the 7x + 5(-12 +7x) = -18

OpenStudy (blackops2luvr):

Not really, i was homeschooled for math, then i was gonna be put in regular school for math, but they had to register me and all that junk, so i was 3 weeks behind, and now were starting this and i have no idea whats going on

OpenStudy (anonymous):

lol well do you know PEMDAS?

OpenStudy (blackops2luvr):

Nope, yesterday was my first day.

OpenStudy (anonymous):

Well the P is Parenthesis

OpenStudy (anonymous):

() that goes first so you do whats in Parenthesis first by first starting off by 5 times -12

OpenStudy (anonymous):

@Legends

OpenStudy (anonymous):

please help? what is an example for a system of non-linear equations for algebra 2? BUT THE EQUATION CANT HAVE WHOLE NUMBER ANSWERS

OpenStudy (anonymous):

hold on @mabitrix

OpenStudy (anonymous):

black do you know what 5 times -12 is ?

OpenStudy (blackops2luvr):

yes lol

OpenStudy (anonymous):

what is it

OpenStudy (blackops2luvr):

-60

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So it now that we've done that whats whats 5 times -7x

OpenStudy (blackops2luvr):

-35

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-35x bring down are x

OpenStudy (anonymous):

Now were done with parenthesis

OpenStudy (anonymous):

we how 7x + -60 -35x = -18

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

Do you know how to put are x's together

OpenStudy (blackops2luvr):

No

OpenStudy (anonymous):

you see the numbers with the x

OpenStudy (anonymous):

you put those together and you put the ones with no x's together

OpenStudy (anonymous):

7x - 35x= -18 +60

OpenStudy (anonymous):

do you know how to do that

OpenStudy (blackops2luvr):

hold on

OpenStudy (anonymous):

k

OpenStudy (blackops2luvr):

No.

OpenStudy (anonymous):

whats 7 -35

OpenStudy (blackops2luvr):

-28

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

let me just tell you the answer hold on

OpenStudy (anonymous):

whats the answer choices

OpenStudy (anonymous):

i got the answer

OpenStudy (anonymous):

@blackops2luvr

OpenStudy (anonymous):

Whats the answers they gave you

OpenStudy (blackops2luvr):

no answer choises